OpenStudy (anonymous):

What is the simplified form of x^2 - 16 / x + 4

jimthompson5910 (jim_thompson5910):

Hint: x^2 - 16 x^2 - 4^2 (x-4)(x+4)

OpenStudy (anonymous):

x - 4, with the restriction x ≠ 4 x + 4, with the restriction x ≠ - 4 x - 4, with the restriction x ≠ - 4 x + 4, with the restriction x ≠ 4

jimthompson5910 (jim_thompson5910):

I'm using the difference of squares rule

OpenStudy (anonymous):

@jim_thompson5910 is it A?

jimthompson5910 (jim_thompson5910):

it is x-4, but the restriction is not x ≠ 4 since 4 is a perfectly valid input in the original expression

jimthompson5910 (jim_thompson5910):

x+4 = 0 x = -4 is the value that makes the denominator zero, so this is the restricted value

OpenStudy (anonymous):

So it would be C

jimthompson5910 (jim_thompson5910):

yes, it's C

OpenStudy (anonymous):

@jim_thompson5910 can u plz explain me more i'm confused

jimthompson5910 (jim_thompson5910):

which part are you confused about?

jimthompson5910 (jim_thompson5910):

if you want, you can ask the question in a separate post

OpenStudy (anonymous):

all

jimthompson5910 (jim_thompson5910):

\[\Large \frac{x^2 - 16}{x+4}\] \[\Large \frac{x^2 - 4^2}{x+4}\] \[\Large \frac{(x-4)(x+4)}{x+4}\] \[\Large \frac{(x-4)\cancel{(x+4)}}{\cancel{x+4}}\] \[\Large x-4\] So \[\Large \frac{x^2 - 16}{x+4}\] simplifies to \[\Large x-4\] Keep in mind that x cannot equal -4 in the original expression. So for the two expressions to be completely equivalent, x cannot equal -4 in the final expression.